# FANformer: Improving Large Language Models Through Effective Periodicity Modeling

### v2

@@ Abstract
Periodicity, as one of the most important basic characteristics, lays the foundation for facilitating structured knowledge acquisition and systematic cognitive processes within human learning paradigms. However, the potential flaws of periodicity modeling in Transformer affect the learning efficiency and establishment of underlying principles from data for large language models (LLMs) built upon it. In this paper, we demonstrate that integrating effective periodicity modeling can improve the learning efficiency and performance of LLMs. We introduce FANformer, which adapts Fourier Analysis Network (FAN) into attention mechanism to achieve efficient periodicity modeling, by modifying the feature projection process of attention mechanism. Extensive experimental results on language modeling show that FANformer consistently outperforms Transformer when scaling up model size and training tokens, underscoring its superior learning efficiency. Our pretrained FANformer-1B exhibits marked improvements on downstream tasks compared to open-source LLMs with similar model parameters or training tokens. Moreover, we reveal that FANformer exhibits superior ability to learn and apply rules for reasoning compared to Transformer. The results position FANformer as an effective and promising architecture for advancing LLMs. Our code and pretrained model are available at https://github.com/YihongDong/FANformer.
@@ 3.1 ATF
$\mathbf{X}_{F}=\text{FANLayer}^{\prime}(\mathbf{X})=[\cos(W_{p}\mathbf{X})||\sin(W_{p}\mathbf{X})||(W_{\bar{p}}\mathbf{X}+B_{\bar{p}})],$ (1)
@@
$[\mathbf{Q}_{F},\mathbf{K}_{F},\mathbf{V}_{F}]=\mathbf{X}_{F}[\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}],$ (2)
@@
where $\mathbf{W}_{O}\in\mathbb{R}^{d_{h}\times d_{h}}$ is the learnable parameter of out projection matrix. Note that $\text{ATF}(\mathbf{X})$ is mathematically equivalent to $\text{Attention}(\text{FANLayer}^{\prime}(\mathbf{X}))$ (the detailed derivations are provided in Appendix I). This equivalence enables a simple yet effective implementation of Multi-head ATF as shown in Figure 2, which can seamlessly incorporate various advancements in traditional attention mechanisms, such as FlashAttention (Dao et al., 2022).
@@ 4 Evaluation
We begin with the implementation details of our experiments (Section 4.1), followed by a comprehensive evaluation of FANformer from three distinct perspectives: First, we investigate the scalability of FANformer by examining its performance trends on language modeling tasks with respect to model size and training tokens (Section 4.2). Second, we evaluate the capabilities of the pre-trained FANformer-1B model across multiple downstream tasks (Section 4.3). Third, we conduct an in-depth empirical analysis of FANformer, including ablation study, learning efficiency, reasoning mechanism, representational capacity, and more (Section 4.4). See Appendix A-G for more experiments.
@@ 4.1 Implementation Details
The experiments are conducted on 80 A100 GPUs. We build FANformer upon the foundation of OLMo (Groeneveld et al., 2024), as it provides a solid pretraining framework of LLMs, with the hyperparameter $p$ set to 0.25 by default. For pretraining FANformer-1B, we randomly sample 1T training tokens from OLMo’s training data, i.e., Dolma (Soldaini et al., 2024). For other experiments, we train models on a smaller sample of Dolma, i.e., Dolma v1_6-sample (AllenAI, 2023), with roughly 10B tokens. The detailed pretraining and experimental setups are provided in Appendix J.
@@ 4.4.1 Ablation Study and Variant Analysis
Variants Param. Training Loss $\downarrow$ V2 Eval Loss $\downarrow$ V2 Eval PPL $\downarrow$ V3 Eval Loss $\downarrow$ V3 Eval PPL $\downarrow$ DownStream Avg Acc. $\uparrow$
Transformer 1.0×1.0$\times$1.0 × 2.889 3.33 30.20 3.07 24.28 53.10
Transformer (FFN $\leftarrow$ FAN) 1.0×1.0$\times$1.0 × 2.880 3.31 29.79 3.05 23.96 53.95
[…]
Transformer + ATM 1.0×1.0$\times$1.0 × 2.890 3.33 30.31 3.07 24.36 53.69
Transformer + ATL 1.0×1.0$\times$1.0 × 2.882 3.31 29.68 3.05 23.97 53.46
FANformer (original FAN) 1.0×1.0$\times$1.0 × 2.893 3.34 30.64 3.07 24.50 53.61
FANformer 1.0×1.0$\times$1.0 × 2.863 3.30 29.40 3.04 23.62 55.19
[…]
Transformer + ATM 1.06×1.06$\times$1.06 × 2.886 3.33 30.18 3.06 24.28 52.86
Transformer + ATL 1.06×1.06$\times$1.06 × 2.879 3.31 29.76 3.05 23.94 54.23
FANformer (original FAN) 1.04×1.04$\times$1.04 × 2.887 3.34 30.57 3.07 24.39 53.13
FANformer 1.04×1.04$\times$1.04 × 2.856 3.29 29.22 3.03 23.47 54.88
@@ Results.
From Table 4.4.1, we have the following findings: 1) FANformer consistently outperforms other variant architectures in both scenarios of the same parameter and same dimension on all evaluation metrics. 2) The performance of Transformer+ATM and Transformer+ATL is notably inferior to that of FANformer, indicating that the core improvement stems from the ATF module we designed. 3) Although Transformer (FFN $\leftarrow$ FAN) yields some improvement, this enhancement is inferior to the gains achieved by FANformer, suggesting that integrating periodicity modeling within attention is more advantageous than FFN on language modeling. 4) Incorporating activation functions such as GELU into the attention mechanism tends to degrade model performance. Specifically, FANformer (original FAN) and Transformer+ATM exhibit weaker performance compared to FANformer and Transformer+ATL, likely because these activation functions suppress certain features, thereby hindering subsequent attention operations.
@@ 5 Related Work
Recent advancements in Transformer architecture focus on overcoming two main limitations: computational inefficiency in long-context processing and limited expressiveness of attention mechanisms. To address long-context inefficiencies, techniques such as sparsity (Child et al., 2019) and local attention (Beltagy et al., 2020), along with innovations in query mechanisms (Shazeer, 2019; Ainslie et al., 2023) and inference acceleration (DeepSeek-AI et al., 2024b), have been proposed. Hardware-level optimizations like FlashAttention (Dao et al., 2022) further reduce GPU memory access overhead. To enhance the expressiveness of attention mechanisms, methods like Probabilistic Attention Keys (Nguyen et al., 2022) and Selective Attention (Leviathan et al., 2024) improve semantic relationship capture and refine attention by suppressing irrelevant features. Additionally, approaches like Differential Transformer address attention noise in long contexts (Ye et al., 2024). Different from previous work, we improve language modeling by addressing the challenge of periodicity modeling in Transformers, which can seamlessly incorporate the aforementioned works for revising the attention mechanism, as demonstrated in the derivation provided in Appendix I.
@@ Appendix B Preliminary Knowledge
[…]
Appendix C Effect of hyperparameter $p$
We systematically investigate the impact of hyperparameter $p$, which controls the proportion of periodicity modeling in FANformer, on model performance across its value range. The experimental results from the 1B-scale FANformer (as shown in Figure 8) demonstrate that our model exhibits strong robustness in terms of training loss and downstream task accuracy, with relatively small performance fluctuations. Furthermore, regardless of the variation in p values, FANformer consistently outperforms the standard Transformer (horizontal baseline). Analysis of experimental results from models of different scales (300M, 1B, 3B) (as shown in Figure 9) reveals a clear trend: larger models tend to exhibit higher optimal p values. This observation suggests that more powerful FANformers are better equipped to extract more intricate latent periodicity features.
@@ Appendix C Effect of hyperparameter $p$
Appendix D Instruction Following with SFT
@@ D.2 Evaluation Setup
Appendix E Computational Cost Analysis
@@
Appendix F Detailed Results of Ablation Study for Section 4.4.1
In ablation study, we report the average results across various tasks on V2 Validation Sets, V3 Validation Sets, and Downstream tasks, with the specific tasks detailed in Section J.3. The complete results are detailed in Table 6 and Table 7.
@@
Appendix G Extended results of Section 4.4.3
@@
Appendix H Case for Section 4.4.4
@@
Appendix I Formal Proof of $\text{ATF}(\mathbf{X})=\text{Attention}(\text{FANLayer}^{\prime}(\mathbf{X}))$
@@
Appendix J Comprehensive Experimental Details
J.1 Detailed training settings of FANformer
@@
J.2 Detailed Setup for Section 4.2
@@
J.3 Validation Set And Downstream Tasks
@@
J.4 Detailed Setup of Case-based and Rule-based Reasoning.
@@
  • Modular addition. The input to the model is “$a+b=$”, and the output is $c$, where $c=(a+b)\mod P$. The values of $a$ and $b$ range from 0 to 112. The constant $P$ is $113$ here.
  • Linear regression. This task involves the model learning a linear regression function. The input is given by “$(a,b)=$”, and the output is $c$, where $c=m\cdot a+n\cdot b+p$. The values of $a$ and $b$ range from 0 to 99. The constants are set as $m=1$, $n=2$, and $p=3$.
@@
J.5 Assessing LLMs’ Proficiency in Capturing Inferential Rules
@@
If Premise, then Conclusion__\__original. True
If Premise, then Conclusion__\__flipped. False
Appendix K Limitations
@@
Second, our work is orthogonal to the existing approaches for revising the attention mechanism, i.e., our work can seamlessly incorporate them, as verified in the derivation provided in Appendix I. There are numerous variants of attention mechanisms, as discussed in the related work (Section 5), such as Flash Attention [Dao et al., 2022], MQA [Shazeer, 2019], and MLA [DeepSeek-AI et al., 2024b]. In this work, we only incorporate Flash Attention for necessary acceleration, while leaving the exploration of other approaches for future work.

### v4

# Reasoning is Periodicity? Improving Large Language Models Through Effective Periodicity Modeling

@@ Abstract
Periodicity, as one of the most important basic characteristics, lays the foundation for facilitating structured knowledge acquisition and systematic cognitive processes within human learning paradigms. However, the potential flaws of periodicity modeling in Transformer affect the learning efficiency and establishment of underlying principles from data for large language models (LLMs) built upon it. In this paper, we demonstrate that integrating effective periodicity modeling can improve the learning efficiency and performance of LLMs. We introduce FANformer, which adapts Fourier Analysis Network (FAN) into attention mechanism to achieve efficient periodicity modeling, by modifying the feature projection process of attention mechanism. Extensive experimental results on language modeling show that FANformer consistently outperforms Transformer when scaling up model size and training tokens, underscoring its superior learning efficiency. Our pretrained FANformer-1B exhibits marked improvements on downstream tasks compared to open-source LLMs with similar model parameters or training tokens. Moreover, we reveal that FANformer exhibits superior ability to learn and apply rules for reasoning compared to Transformer. The results position FANformer as an effective and promising architecture for advancing LLMs. Our code is available at https://github.com/YihongDong/FANformer..
@@ 3.1 ATF
[…]
@@
$\mathbf{Q}_{F},\mathbf{K}_{F},\mathbf{V}_{F}]=\mathbf{X}_{F}[\mathbf{W}_{Q},\mathbf{W}_{K},\mathbf{W}_{V}],$ (2)
@@
where $\mathbf{W}_{O}\in\mathbb{R}^{d_{h}\times d_{h}}$ is the learnable parameter of out projection matrix. Note that $\text{ATF}(\mathbf{X})$ is mathematically equivalent to $\text{Attention}(\text{FANLayer}^{\prime}(\mathbf{X}))$ (the detailed derivations are provided in Appendix L). This equivalence enables a simple yet effective implementation of Multi-head ATF as shown in Figure 2, which can seamlessly incorporate various advancements in traditional attention mechanisms, such as FlashAttention (Dao et al., 2022).
@@ 4 Evaluation
We begin with the implementation details of our experiments (Section 4.1), followed by a comprehensive evaluation of FANformer from three distinct perspectives: First, we investigate the scalability of FANformer by examining its performance trends on language modeling tasks with respect to model size and training tokens (Section 4.2). Second, we evaluate the capabilities of the pre-trained FANformer-1B model across multiple downstream tasks (Section 4.3). Third, we conduct an in-depth empirical analysis of FANformer, including ablation study, learning efficiency, reasoning mechanism, representational capacity, and more (Section 4.4). See Appendix A-J for more experiments.
@@ 4.1 Implementation Details
The experiments are conducted on 80 A100 GPUs. We build FANformer upon the foundation of OLMo (Groeneveld et al., 2024), as it provides a solid pretraining framework of LLMs, with the hyperparameter $p$ set to 0.25 by default. For pretraining FANformer-1B, we randomly sample 1T training tokens from OLMo’s training data, i.e., Dolma (Soldaini et al., 2024). For other experiments, we train models on a smaller sample of Dolma, i.e., Dolma v1_6-sample (AllenAI, 2023), with roughly 10B tokens. The detailed pretraining and experimental setups are provided in Appendix M.
@@ 4.4.1 Ablation Study and Variant Analysis
Variants Param.
Training
Loss $\downarrow$
V2 Eval
Loss $\downarrow$
V2 Eval
PPL $\downarrow$
V3 Eval
Loss $\downarrow$
V3 Eval
PPL $\downarrow$
DownStream
Avg Acc. $\uparrow$
Transformer $1.0\times$ 2.889 3.33 30.20 3.07 24.28 53.10
Transformer (FFN $\leftarrow$ FAN) $1.0\times$ 2.880 3.31 29.79 3.05 23.96 53.95
Same Parameter
Transformer + ATM $1.0\times$ 2.890 3.33 30.31 3.07 24.36 53.69
Transformer + ATL $1.0\times$ 2.882 3.31 29.68 3.05 23.97 53.46
FANformer (original FAN) $1.0\times$ 2.893 3.34 30.64 3.07 24.50 53.61
FANformer $1.0\times$ 2.863 3.30 29.40 3.04 23.62 55.19
Same Dimension
Transformer + ATM $1.06\times$ 2.886 3.33 30.18 3.06 24.28 52.86
Transformer + ATL $1.06\times$ 2.879 3.31 29.76 3.05 23.94 54.23
FANformer (original FAN) $1.04\times$ 2.887 3.34 30.57 3.07 24.39 53.13
FANformer $1.04\times$ 2.856 3.29 29.22 3.03 23.47 54.88
@@ Results.
From Table 2, we have the following findings: 1) FANformer consistently outperforms other variant architectures in both scenarios of the same parameter and same dimension on all evaluation metrics. 2) The performance of Transformer+ATM and Transformer+ATL is notably inferior to that of FANformer, indicating that the core improvement stems from the ATF module we designed. 3) Although Transformer (FFN $\leftarrow$ FAN) yields some improvement, this enhancement is inferior to the gains achieved by FANformer, suggesting that integrating periodicity modeling within attention is more advantageous than FFN on language modeling. 4) Incorporating activation functions such as GELU into the attention mechanism tends to degrade model performance. Specifically, FANformer (original FAN) and Transformer+ATM exhibit weaker performance compared to FANformer and Transformer+ATL, likely because these activation functions suppress certain features, thereby hindering subsequent attention operations.
@@ 5 Related Work
Recent advancements in Transformer architecture focus on overcoming two main limitations: computational inefficiency in long-context processing and limited expressiveness of attention mechanisms. To address long-context inefficiencies, techniques such as sparsity (Child et al., 2019) and local attention (Beltagy et al., 2020), along with innovations in query mechanisms (Shazeer, 2019; Ainslie et al., 2023) and inference acceleration (DeepSeek-AI et al., 2024b), have been proposed. Hardware-level optimizations like FlashAttention (Dao et al., 2022) further reduce GPU memory access overhead. To enhance the expressiveness of attention mechanisms, methods like Probabilistic Attention Keys (Nguyen et al., 2022) and Selective Attention (Leviathan et al., 2024) improve semantic relationship capture and refine attention by suppressing irrelevant features. Additionally, approaches like Differential Transformer address attention noise in long contexts (Ye et al., 2024). Different from previous work, we improve language modeling by addressing the challenge of periodicity modeling in Transformers, which can seamlessly incorporate the aforementioned works for revising the attention mechanism, as demonstrated in the derivation provided in Appendix L.
@@ 6 Conclusion
7 Acknowledgement
This research is supported by the National Key R&\&D Program under Grant No. 2023YFB4503801, the National Natural Science Foundation of China under Grant No. 62192733, 62192730, 62192731, the Major Program (JD) of Hubei Province (No.2023BAA024).
@@ Appendix B Preliminary Knowledge
where $W_{p}$ and $W_{\bar{p}}$ are learnable projection matrices, $B_{\bar{p}}$ is a bias term, $\sigma$ denotes an activation function, and $\|$ represents concatenation. Compared to MLP layer, FAN layer explicitly encodes periodic patterns through Fourier series while maintaining general-purpose modeling capabilities.
Appendix C Experiments on Code Generation and Math tasks
We conduct experiments on code generation tasks (i.e., HumanEval and MBPP) and human-written math tasks (i.e., GSM8K) compared to our baseline, OLMo [Groeneveld et al., 2024]. The results show that our FANformer achieves clear and consistent improvements compared with OLMo on all three benchmarks.
Table 4: Comparison of LLMs on coding and math benchmarks.
LLMs Training Tokens HumanEval MBPP GSM8K
OLMo-1B 3T checkpoint 5.2 3.1 8.9
FANformer-1B 1T from scratch 6.3 5.4 15.7
Appendix D Effect of hyperparameter $p$
We systematically investigate the impact of hyperparameter $p$, which controls the proportion of periodicity modeling in FANformer, on model performance across its value range. The experimental results from the 1B-scale FANformer (as shown in Figure 8) demonstrate that our model exhibits strong robustness in terms of training loss and downstream task accuracy, with relatively small performance fluctuations. Furthermore, regardless of the variation in p-values, FANformer consistently outperforms the standard Transformer (horizontal baseline). Analysis of experimental results from models of different scales (300M, 1B, 3B) (as shown in Figure 9) reveals a clear trend: larger models tend to exhibit higher optimal p values. This observation suggests that more powerful FANformers are better equipped to extract more intricate latent periodicity features.
@@ Appendix D Effect of hyperparameter $p$
Appendix E Instruction Following with SFT
@@ E.2 Evaluation Setup
Appendix F Computational Cost Analysis
@@
Appendix G Inference Speed and GPU Memory Usage
We conduct experiments on the inference speed and GPU memory usage of FANformer relative to Transformer in deployment settings, and have added the experimental results to our revised manuscript. The results show that it add little latency. The configuration of benchmark test: we run for 20 iterations on a single GPU of A100 80G with a fixed sequence length of 4096 tokens and float16 precision.
Table 7: Inference speed and memory usage comparison (Sequence Length=4096).
Metric OLMo-1B FANformer-1B Difference
Forward Pass Time 141.49 ms 142.88 ms +1.39 ms (+0.98%)
Allocated Memory 4642.69 MB 4738.86 MB +96.17 MB (+2.1%)
Peak Memory 6610.70 MB 6706.88 MB +96.18 MB (+1.5%)
Appendix H Different with Mamba
First, our approach is fundamentally different from SSMs. SSMs model periodicity along the sequence dimension, while our FANformer models it along the feature dimension. Second, our motivation is distinct from that of Mamba. Mamba [Gu and Dao, 2023] is primarily developed to overcome the quadratic computational complexity of Transformers and improve inference efficiency, while our approach is designed to improve learning efficiency and performance of Transformers through effective periodicity modeling.
Appendix I Detailed Results of Ablation Study for Section 4.4.1
In ablation study, we report the average results across various tasks on V2 Validation Sets, V3 Validation Sets, and Downstream tasks, with the specific tasks detailed in Section M.3. The complete results are detailed in Table 8 and Table 9.
@@ Appendix H Different with Mamba
Appendix J Extended results of Section 4.4.3
@@
Appendix K Case for Section 4.4.4
@@
Appendix L Formal Proof of $\text{ATF}(\mathbf{X})=\text{Attention}(\text{FANLayer}^{\prime}(\mathbf{X}))$
@@
Appendix M Comprehensive Experimental Details
M.1 Detailed training settings of FANformer
@@
M.2 Detailed Setup for Section 4.2
@@
M.3 Validation Set And Downstream Tasks
@@
M.4 Detailed Setup of Case-based and Rule-based Reasoning.
@@
  • Modular addition. The input to the model is “a+b=a+b=”, and the output is $c$, where $c=(a+b)\mod P$. The values of $a$ and $b$ range from 0 to 112. The constant $P$ is $113$ here.
  • Linear regression. This task involves the model learning a linear regression function. The input is given by “(a,b)=(a,b)=”, and the output is $c$, where $c=m\cdot a+n\cdot b+p$. The values of $a$ and $b$ range from 0 to 99. The constants are set as $m=1$, $n=2$, and $p=3$.
@@
M.5 Assessing LLMs’ Proficiency in Capturing Inferential Rules
@@
If Premise, then Conclusion_\_original. True
If Premise, then Conclusion_\_flipped. False
Appendix N Limitations
@@
Second, our work is orthogonal to the existing approaches for revising the attention mechanism, i.e., our work can seamlessly incorporate them, as verified in the derivation provided in Appendix L. There are numerous variants of attention mechanisms, as discussed in the related work (Section 5), such as Flash Attention [Dao et al., 2022], MQA [Shazeer, 2019], and MLA [DeepSeek-AI et al., 2024b]. In this work, we only incorporate Flash Attention for necessary acceleration, while leaving the exploration of other approaches for future work.